# Core-Set Selection for Data-efficient Land Cover Segmentation

### v1

@@ Abstract
The increasing accessibility of remotely sensed data and the potential of such data to inform large-scale decision-making has driven the development of deep learning models for many Earth Observation tasks. Traditionally, such models must be trained on large datasets. However, the common assumption that broadly larger datasets lead to better outcomes tends to overlook the complexities of the data distribution, the potential for introducing biases and noise, and the computational resources required for processing and storing vast datasets. Therefore, effective solutions should consider both the quantity and quality of data. In this paper, we propose six novel core-set selection methods for selecting important subsets of samples from remote sensing image segmentation datasets that rely on imagery only, labels only, and a combination of each. We benchmark these approaches against a random-selection baseline on three commonly used land cover classification datasets: DFC2022, Vaihingen, and Potsdam. In each of the datasets, we demonstrate that training on a subset of samples outperforms the random baseline, and some approaches outperform training on all available data. This result shows the importance and potential of data-centric learning for the remote sensing domain. The code is available at https://github.com/keillernogueira/data-centric-rs-classification/.
@@ I Introduction
To this end, core-set selection is a technique that focuses on finding and using the most valuable samples for training deep learning models, thus preserving the essential characteristics and insights of the entire dataset while maintaining or even enhancing overall model performance. Such a paradigm can assist in several aspects such as improved computational efficiency, cost-effective data handling, enhanced model performance, effective use of labeled data, or efficient labeling of unlabeled data [5, 6, 7, 8].
@@
Given its importance, several papers have addressed core-set selection, usually employing methods such as clustering algorithms and gradient approximation [9, 10, 11]. Some works perform the core-set selection before training the final machine-learning model [9, 10] whereas others perform a new selection at each training epoch [11, 12]. The former is agnostic regarding the model that is finally applied to the established training data but the selection process is performed only once. on the other hand, the latter is more optimized for the specific machine learning model, but it requires more computational resources since core-set selection is performed multiple times during training. Since we aim to establish a general baseline for core-set selection methods for remote sensing data, we focus on approaches that are independent of the used downstream model. Finally, although some of these works perform core-set selection for image classification, to the best of our knowledge, there have been no initiatives exploiting this paradigm for remote sensing image segmentation, which is the focus of our work.
@@
In this paper, we propose novel core-set selection approaches for remote sensing image segmentation based on several distinct premises, which rely on imagery only, labels only, and a combination of each. Specifically, given the training samples (images and their segmentation masks), the proposed approaches rank these examples from most to least valuable for training. This ranking is based on how representative each sample is, according to specific criteria. Instead of training the deep learning-based models using all available samples, we leverage the aforementioned ranking to select the most representative examples (core-set) and train the models accordingly. By doing so, we can not only reduce the training time but also improve the overall effectiveness by filtering out non-representative and/or noisy samples. In practice, the main contributions of this paper are the following:
@@
  • Novel core-set selection approaches for remote sensing image segmentation, each based on different premises; and
  • A ranking mechanism to evaluate the representativeness of training samples, enabling the selection and adaptation of the most valuable examples depending on the problem.
@@
Overall, this work, an outcome of the Data-Centric Land Cover Classification Challenge part of the Workshop on Machine Vision for Earth Observation and Environment Monitoring (MVEO) 2024, fills a critical gap in the literature and demonstrates the potential of core-set selection in advancing remote sensing image segmentation as well as data creation and labeling.
@@
Figure 1: General overview of core-set selection. An input data set is first processed by a core-set selection algorithm that, based on some criteria, prioritizes certain samples over others (represented by the size of the blue circles). Based on this, it is possible to select the core-set data depending on the amount of data one would like to retain (illustrated by red, orange, and green dotted circles). Finally, the selected core-set is used to train a machine learning model, thus reducing the training time while maintaining, or even improving, task performance.
@@ II Related work
Another reason is to improve model performance by enhancing the learning process and reducing overfitting through filtering out noisy or incorrect data points, thus creating cleaner datasets [13]. In remote sensing, predominantly prior knowledge, label information or an existing model is used to identify a clean core-set [14, 15]. Santos et al. [10], for example, use clustering for satellite time series to identify samples that are mislabeled or have low accuracy, with the goal of removing them from the training set to avoid a decrease in model performance. Moreover, many methods have been developed for data with known sources of uncertainty, such as clouds [16, 17, 18]. Furthermore, this reason is directly related to the enhancement of the accuracy and robustness of machine learning models by removing low-quality or redundant samples. Such models can perform as well as, or even better than, those trained on the full dataset (e.g., [19]).
Another reason for core-set selection is to support clear and non-misleading explanations of a model and the data. Generally, with the field of explainable machine learning, new methods are introduced to calculate importance scores, sensitivities, or contributions of features and interpret them as relevance [20]. However, redundancies and correlations distort the derived insights, therefore they should be removed before interpreting and explaining the results. With the goal to analyze geospatial air quality estimations and the relevance of specific measurement locations, Stadtler et al. [21] demonstrated that removing redundant samples only slightly decreases test accuracy, as these samples are not relevant for training.
In general, the principles of core-set selection are closely related to areas like active learning and self-training. For unlabeled datasets, core-set selection can guide efficient labeling by identifying the most representative samples. This optimizes resource allocation for manual annotation. In case data is already labeled, core-set selection can help prioritize the most informative samples. This maximizes the use of labeled data and may reduce the need for further labeling efforts.
Overall, although most of the aforementioned works perform core-set selection for image classification, to the best of our knowledge, our work is the first research to design core-set selection techniques specifically for remote sensing image segmentation.
@@ III Core-set selection methods
Given a set of satellite images $\mathbf{X}$ and their corresponding label masks $\mathbf{M}$ (also known as segmentation maps), we propose several core-set selection methods that assign an importance score, $s_{i}$, to the $i$-th sample $(X_{i},M_{i})$. These scores range from 0 (least valuable for training) to 1 (most valuable for training). The aim of these methods is to rank the samples based on their informativeness, allowing us to select a subset of the data — a core-set — that can achieve good model performance with reduced training time and data size.
The proposed methods are categorized into: label-based methods, which only rely on the label masks $\mathbf{M}$, image-based methods, which use information only from the input images $\mathbf{X}$, and combined methods that integrate both sources.
For a given training budget $b$, the core-set consists of the top-$b$ samples ranked by their scores. Next, we describe the methods in detail.
@@ III-A Label Complexity (LC)
LC is a label-based method that scores a sample based on the complexity of its label mask $M_{i}$. The underlying assumption is that samples with high-complexity label masks are more informative for training segmentation models. The complexity of a label mask is quantified using entropy, which measures the diversity of the class distribution in the label mask. For each sample $i$, we compute the score $s_{i}^{\text{LC}}$ based on the entropy $H(M_{i})$ as
$s_{i}^{\text{LC}}=1-H(M_{i})=1+\sum_{c=1}^{C}p_{i,c}\log_{C}(p_{i,c}),$ (1)
@@
Higher scores correspond to samples that have a more uniform distribution of class labels with potentially more informative label masks, while low scores correspond to samples that are dominated by a single class.
@@ III-B Feature Space Diversity (FD)
The FD method is an image-based approach that aims to select a diverse core-set of samples from $\mathbf{X}$ based on each sample’s embedding from a pre-trained deep learning model.
@@
Next, we group the feature embeddings into $K$ clusters using the K-Means algorithm. We search for a value of $K$ that minimizes the average Vendi score [24] across clusters. The Vendi score is a measure of diversity over a set of vectors – by minimizing the average within-cluster diversity we ensure that a sample from that cluster is representative of the others. Starting with $K=2$, we cluster all image embeddings with K-Means, measure the average per-cluster Vendi score, then increment $K$, and repeat until the change in the average Vendi score falls below a threshold of $\delta$ (we choose $\delta=0.5\%$) for three iterations.
Given a clustering of the samples, we sequentially choose one sample randomly from within each cluster in a round-robin fashion (the first cluster is randomly selected from the set of $K$ clusters), resulting in ordered $K$-sized segments of cross-cluster samples. The first sample has the highest importance score $s_{i}^{\text{FD}}=1$, while the last selected sample receives the lowest score $s_{i}^{\text{FD}}=0$.
@@ III-C Complexity/Diversity Hybrid (LC/FD)
The LC/FD method is a hybrid method that combines the most important samples from the LC and FD methods, following an assumption that diversity is important when working with small datasets, while label complexity becomes increasingly important for medium to large datasets.
Specifically, the LC/FD method is a hybrid approach that uses the ranked lists of samples from LC and FD, denoted as $\mathcal{R}_{\text{LC}}$ and $\mathcal{R}_{\text{FD}}$, respectively. A cutoff point $m$ is defined, and the hybrid ranking is constructed by taking the top $m$ samples from $\mathcal{R}_{\text{FD}}$, followed by all samples from $\mathcal{R}_{\text{LC}}$ that are not already included. This ensures that the selected core-set includes a mix of label-complex samples and feature-diverse samples.
@@ III-D Feature Activation (FA)
The FA method is an image-based approach that uses statistics from image embeddings created by a pre-trained neural network to rank samples.
First, a ResNet-18 [22], pre-trained on the ImageNet dataset, is used to extract the image embeddings (after the final channel-wise pooling layer), resulting in a feature vector, $F_{i}\in\mathbb{R}^{512}$, that encodes higher-level semantic information per image.
By construction, all values in $F_{i}$ are non-negative (via application of a ReLU function within the network). We assume that samples with high activation magnitudes (large mean) and significant variations across different dimensions (high standard deviation) in feature space are likely to carry more relevant information. Therefore, we compute the score $s_{i}^{\text{FA}}$ with a combination of the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ (both scaled to $(0-1]$) of the embedding vector $F_{i}$ as
@@
According to the aforementioned assumption, samples with low scores are likely to be less important for training, may have lower diversity, or contain more noise, among other issues.
@@ III-E Class Balance (CB)
Similar to the LC method, the CB method is a label-based method that aims to select a subset of samples with a uniform class distribution, but using a non-greedy approach.
The method consists of $N$ steps, where $N$ refers to the number of samples in the dataset. In each step, the most suitable sample is selected from the dataset to ensure that the overall class distribution of the selected samples approaches a uniform distribution. Specifically, the most suitable sample is the sample that maximizes the entropy of the class distribution of the union of the current core-set with the selected sample.
The order in which the samples are selected determines their importance score: the first sample is ranked as the most important and the last sample is ranked as least important. Formally, let $r_{i}$ be the rank of sample $i$, then
@@ III-F Feature Activation/Class Balance Hybrid (FA/CB)
The FA/CB method is a hybrid method that uses a weighted ensemble of importance scores calculated by the previous two methods, i.e.
@@ IV-A1 IEEE GRSS Data Fusion Contest 2022 (DFC2022) Dataset
The DFC2022 dataset [25] was released as part of the annual IEEE GRSS Data Fusion Contest.
This dataset consists of images gathered in and around 19 urban areas from different regions in France. Each sample of this dataset contains a high-resolution RGB aerial image and its corresponding label mask, both having approximately $2000\times 2000$ pixels and a spatial resolution of 50 cm per pixel, along with a Digital Elevation Model (DEM) with $1000\times 1000$ pixels at a spatial resolution of 100cm/pixel. For our experiments, we resample the DEM data to match the dimensions of the image and label mask. The masks in this dataset contain 14 classes: “urban fabric”, “industrial, commercial, public, military, private and transport units”, “mine, dump and construction sites”, “artificial non-agricultural vegetated areas”, “arable land (annual crops)”, “permanent crops”, “pastures”, “complex and mixed cultivation patterns”, “orchards at the fringe of urban classes”, “forests”, “herbaceous vegetation associations”, “open spaces with little or no vegetation”, “wetlands”, and “water”. Samples of this dataset are shown in Figure 2.
@@
Figure 2: Some images (first row) of the DFC2022 dataset [25], their DEM data (second row), and the respective reference data (third row).
@@ IV-A2 ISPRS Vaihingen and Potsdam Datasets
The Vaihingen and Potsdam datasets [26] were released for the 2D semantic labeling contest of the International Society for Photogrammetry and Remote Sensing (ISPRS). Both datasets consist of aerial imagery, Digital Surface Model (DSM) data, and label masks, as shown in Figure 3. The Vaihingen dataset contains 33 patches, with an average size of $2494\times 2064$ pixels. The aerial images have three bands (near-infrared, red, and green) with a spatial resolution of 9 cm per pixel. The Postdam dataset contains 38 tiles of $6000\times 6000$ pixels. The imagery consists of four bands (red, green, blue, and near-infrared) with a spatial resolution of 5 cm per pixel. The label masks in both datasets contain six classes: “impervious surfaces”, “building”, “low vegetation”, “tree”, “car”, “clutter/background”.
@@
Figure 3: Some images (first row) of the Vaihingen and Potsdam datasets [26], their DSM data (second row), and the respective reference data (third row).

### v2

@@ Abstract
The increasing accessibility of remotely sensed data and the potential of such data to inform large-scale decision-making has driven the development of deep learning models for many Earth Observation tasks. Traditionally, such models must be trained on large datasets. However, the common assumption that broadly larger datasets lead to better outcomes tends to overlook the complexities of the data distribution, the potential for introducing biases and noise, and the computational resources required for processing and storing vast datasets. Therefore, effective solutions should consider both the quantity and quality of data. In this paper, we introduce six basic core-set selection methods for selecting important subsets of examples from remote sensing image segmentation datasets that rely on imagery only, labels only, and a combination of each. We benchmark these approaches against two traditional baselines on three widely used land-cover classification datasets - DFC2022, Vaihingen, and Potsdam - thus establishing a general and comprehensive baseline for future works. In each of the datasets, we demonstrate that the proposed methods outperform the baselines across various settings, with some approaches even selecting core sets that surpass training on all available data. For example, with the DFC2022 dataset, we find a subset of size 10% that, when used in training, results in a slightly better performing model compared to using the entire dataset. This result shows the importance and potential of data-centric learning for the remote sensing domain. The code is available at https://github.com/keillernogueira/data-centric-rs-classification/.
@@ I Introduction
To this end, core-set selection is a technique that focuses on finding and using the most valuable examples for training deep learning models, thus preserving the essential characteristics and insights of the entire dataset while maintaining or even enhancing overall model performance. Such a paradigm can assist in several aspects such as improved computational efficiency, cost-effective data handling, enhanced model performance, effective use of labeled data, or efficient labeling of unlabeled data [5, 6, 7, 8].
@@
Given its importance, several papers have addressed core-set selection, usually employing methods such as clustering algorithms and gradient approximation [9, 10, 11]. Some works perform the core-set selection before training the final machine-learning model [9, 10] whereas others perform a new selection at each training epoch [12, 11, 13]. The former is agnostic regarding the model that is finally applied to the established training data but the selection process is performed only once. On the other hand, the latter is more optimized for the specific machine learning model, but it requires more computational resources since core-set selection is performed multiple times during training. Since we aim to establish a general baseline for core-set selection methods for remote sensing data, we focus on approaches that are independent of the downstream model (i.e., model agnostic), even because these can be easily adapted and used in different scenarios, applications, and tasks, including with modern architectures, such as vision-language and foundation models. Finally, although some of these works perform core-set selection for image classification, to the best of our knowledge, there have been no initiatives exploiting this paradigm for remote sensing image segmentation, which is the focus of our work.
@@
Towards establishing this general and comprehensive baseline, this paper introduces and benchmarks six basic core-set selection approaches for remote sensing image segmentation based on several distinct premises, that rely on imagery only, labels only, and a combination of each. Specifically, given the training instances (i.e., images and their corresponding segmentation masks), the proposed approaches rank the examples from most to least valuable for training, based on the representativeness of each example, according to specific criteria. Instead of training the deep learning-based models using all available data, we leverage the aforementioned ranking to select the most representative examples (core-set) and train the models accordingly. By doing so, we can not only reduce the training time but also improve the overall effectiveness by filtering out non-representative and/or noisy examples. The main contributions of this paper are the following:
@@
  • Introduction of six core-set selection approaches for remote sensing image segmentation, each based on different premises; and
  • A full set of experiments comparing these approaches against two common baselines on three widely used datasets, thus establishing a benchmark for future research in core-set selection.
@@
Overall, this work, an outcome of the Data-Centric Land Cover Classification Challenge of the Workshop on Machine Vision for Earth Observation and Environment Monitoring (MVEO) 2023, fills a critical gap in the literature and demonstrates the potential of core-set selection in advancing remote sensing image segmentation as well as data curation and labeling.
@@
Figure 1: General overview of core-set selection. An input data set is first processed by a core-set selection algorithm that, based on some criteria, prioritizes certain examples over others (represented by the size of the blue circles). Based on this, it is possible to select the core-set data depending on the amount of data one would like to retain (illustrated by red, orange, and green dotted circles). Finally, the selected core-set is used to train a machine learning model, thus reducing the training time while maintaining, or even improving, task performance.
@@ II Related work
Another reason is to improve model performance by enhancing the learning process and reducing overfitting through filtering out noisy or incorrect data points, thus creating cleaner datasets [14]. In remote sensing, predominantly prior knowledge, label information, or an existing model is used to identify a clean core-set [15, 16]. Santos et al. [10], for example, use clustering for satellite time series to identify instances that are mislabeled or have low accuracy, with the goal of removing them from the training set to avoid a decrease in model performance. Moreover, many methods have been developed for data with known sources of uncertainty, such as clouds [17, 18, 19]. Furthermore, this reason is directly related to the enhancement of the accuracy and robustness of machine learning models by removing low-quality or redundant examples. Such models can perform as well as, or even better than, those trained on the full dataset (e.g., [20, 21, 22]).
Another reason for core-set selection is to support clear and non-misleading explanations of a model and the data. Generally, with the field of explainable machine learning, new methods are introduced to calculate importance scores, sensitivities, or contributions of features and interpret them as relevance [23]. However, redundancies and correlations distort the derived insights, therefore they should be removed before interpreting and explaining the results. With the goal to analyze geospatial air quality estimations and the relevance of specific measurement locations, Stadtler et al. [24] demonstrated that removing redundant examples only slightly decreases test accuracy, as these are not relevant for training.
In general, the principles of core-set selection are closely related to areas like active learning [12, 25, 26] and self-training [27]. For unlabeled datasets, core-set selection can guide efficient labeling by identifying the most representative examples. This optimizes resource allocation for manual annotation - a common objective in active learning scenarios [12, 25, 26]. In case the data is already labeled, core-set selection can help prioritize the most informative examples. This maximizes the use of labeled data and may reduce the need for further labeling efforts.
Overall, although most of the aforementioned works perform core-set selection for image classification, to the best of our knowledge, our work is the first research to design and benchmark core-set selection techniques specifically for remote sensing image segmentation.
@@ III Core-set selection methods
Given a set of satellite images $\mathbf{X}$ and their corresponding label masks $\mathbf{M}$ (also known as segmentation maps), we introduce six core-set selection methods that assign an importance score, $s_{i}$, to the $i$-th instance $(X_{i},M_{i})$. These scores range from 0 (least valuable for training) to 1 (most valuable for training). The aim of these methods is to rank the examples based on their informativeness, allowing us to select a subset of the data — a core-set — that can achieve good model performance with reduced training time and data size.
The proposed methods are categorized into: label-based methods, which only rely on the label masks $\mathbf{M}$, image-based methods, which use the information only from the input images $\mathbf{X}$, and combined methods that integrate both sources.
[…]
@@ III-A Label Complexity (LC)
LC is a label-based method that scores an instance based on the complexity of its label mask $M_{i}$. The underlying assumption is that examples with high-complexity label masks are more informative for training segmentation models. It is important to emphasize that this approach does not explicitly guarantee representativeness but instead prioritizes complex label masks to potentially generate more informative training signals, which may lead to improved model performance.
The complexity of a label mask is quantified using the entropy of the class distribution in the label mask – high entropy class distributions will include more and mixed classes, while low or zero entropy class distributions will be dominated by a single class. Precisely, for each instance $i$, we compute the score $s_{i}^{\text{LC}}$ based on the entropy $H(M_{i})$ as:
$s_{i}^{\text{LC}}=H(M_{i})=-\sum_{c=1}^{C}p_{i,c}\log_{C}(p_{i,c}),$ (1)
@@
Higher scores correspond to examples that have a more uniform distribution of class labels with potentially more informative label masks, while low scores correspond to examples that are dominated by a single class.
@@ III-B Feature Space Diversity (FD)
The FD method is an image-based approach that aims to select a diverse core-set of examples from $\mathbf{X}$ based on each example’s embedding from a pre-trained deep learning model.
@@
Next, we group the feature embeddings into $K$ clusters using the K-Means algorithm. We search for a value of $K$ that minimizes the average Vendi score [30] across clusters. The Vendi score is a measure of diversity over a set of vectors – by minimizing the average within-cluster diversity we ensure that an instance from that cluster is representative of the others. Starting with $K=2$, we cluster all image embeddings with K-Means, measure the average per-cluster Vendi score, then increment $K$, and repeat until the change in the average Vendi score falls below a threshold of $\delta$ (we choose $\delta=0.5\%$) for three iterations.
Given a clustering of the examples, we sequentially choose one instance randomly from within each cluster in a round-robin fashion (the first cluster is randomly selected from the set of $K$ clusters), resulting in ordered $K$-sized segments of cross-cluster examples. The first example has the highest importance score $s_{i}^{\text{FD}}=1$, while the last selected example receives the lowest score $s_{i}^{\text{FD}}=0$.
It is important to highlight here that this random selection step does not inherently produce a fixed ordering of scores. However, this does not impact the primary objective of this method, which is to ensure diversity among the selected examples.
@@ III-C Complexity/Diversity Hybrid (LC/FD)
The LC/FD method is a hybrid method that combines the most important examples from the LC and FD methods, following an assumption that diversity is important when working with small datasets, while label complexity becomes increasingly important for medium to large datasets.
Specifically, the LC/FD method is a hybrid approach that uses the ranked lists of examples from LC and FD, denoted as $\mathcal{R}_{\text{LC}}$ and $\mathcal{R}_{\text{FD}}$, respectively. A cutoff point $m$ is defined, and the hybrid ranking is constructed by taking the top $m$ examples from $\mathcal{R}_{\text{FD}}$, followed by all examples from $\mathcal{R}_{\text{LC}}$ that are not already included. This ensures that the selected core-set includes a mix of label-complex examples and feature-diverse instances.
@@ III-D Feature Activation (FA)
The FA method is an image-based approach that uses statistics from image embeddings created by a pre-trained neural network to rank examples.
First, a ResNet-18 [28], pre-trained on the ImageNet dataset, is used to extract the image embeddings (after the final spatial pooling layer), resulting in a feature vector, $F_{i}\in\mathbb{R}^{512}$, that encodes higher-level semantic information per image.
By construction, all values in $F_{i}$ are non-negative (via application of a ReLU function within the network). We assume that examples with high activation magnitudes (large mean) and significant variations across different dimensions (high standard deviation) in feature space are likely to carry more relevant information. Therefore, we compute the score $s_{i}^{\text{FA}}$ with a combination of the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ (both scaled to (01](0-1]( 0 - 1 ]) of the embedding vector $F_{i}$ as:
@@
According to the aforementioned assumption, examples with low scores are likely to have lower diversity, contain more noise, etc, and are therefore likely to be less important for training.
@@ III-E Class Balance (CB)
Similar to the LC, the CB method is a label-based technique that aims to select a subset of examples with a uniform class distribution by using a time-efficient strategy that preprocesses and computes the class distribution of each image for subset selection.
The method consists of $N$ steps, where $N$ refers to the number of examples in the dataset. In each step, the most suitable instance is selected from the dataset to ensure that the overall class distribution of the selected examples approaches a uniform distribution. Specifically, the most suitable instance is the instance that maximizes the entropy of the class distribution of the union of the current core-set with the selected example.
The order in which the examples are selected determines their importance score: the first instance is ranked as the most important and the last example is ranked as least important. Formally, let $r_{i}$ be the rank of instance $i$, then:
@@ III-F Feature Activation/Class Balance Hybrid (FA/CB)
The FA/CB method is a hybrid method that uses a weighted ensemble of importance scores calculated by the previous two methods, that is:
@@ IV-A1 IEEE GRSS Data Fusion Contest 2022 (DFC2022) Dataset
The DFC2022 dataset [31] was released as part of the annual IEEE GRSS Data Fusion Contest. This dataset consists of images gathered in and around 19 urban areas from different regions in France. Each instance of this dataset contains a high-resolution RGB aerial image and its corresponding segmentation mask, both having approximately $\times 2000$ 2000 pixels and a spatial resolution of 50 cm per pixel, along with a Digital Elevation Model (DEM) with $\times 1000$ 1000 pixels at a spatial resolution of 100cm/pixel. For our experiments, we resample the DEM data to match the dimensions of the image and mask. The masks in this dataset contain 14 classes: “urban fabric”, “industrial, commercial, public, military, private and transport units”, “mine, dump and construction sites”, “artificial non-agricultural vegetated areas”, “arable land”, “permanent crops”, “pastures”, “complex and mixed cultivation patterns”, “orchards at the fringe of urban classes”, “forests”, “herbaceous vegetation associations”, “open spaces with little or no vegetation”, “wetlands”, and “water”. Examples of this dataset are shown in Figure 2.
@@
Figure 2: Example images (first row) of the DFC2022 dataset [31], their DEM data (second row), and the respective reference data (third row).
@@ IV-A2 ISPRS Vaihingen and Potsdam Datasets
The Vaihingen and Potsdam datasets [32] were released for the 2D semantic labeling contest of the International Society for Photogrammetry and Remote Sensing (ISPRS). Both datasets consist of aerial imagery, Digital Surface Model (DSM) data, and label masks, as shown in Figure 3. The Vaihingen dataset contains 33 patches, with an average size of 2494×20642494$\times$ 20642494 × 2064 pixels. The aerial images have three bands (near-infrared, red, and green) with a spatial resolution of 9 cm per pixel. The Postdam dataset contains 38 tiles of $\times 6000$ 6000 pixels. The imagery consists of four bands (red, green, blue, and near-infrared) with a spatial resolution of 5 cm per pixel. The label masks in both datasets contain six classes: “impervious surfaces”, “building”, “low vegetation”, “tree”, “car”, and “clutter/background”.
@@
Figure 3: Example images (first row) of the Vaihingen and Potsdam datasets [32], their DSM data (second row), and the respective reference data (third row).